OpenStudy (anonymous):

f(x)=(x^2)/(x-9)^2 Find the vertical and horizontal asymptotes. My problem is I don't know the way to find the vertical and horizontal asymptotes. Is there any algebraic way to find them?

OpenStudy (shubhamsrg):

find that x for which y tends to infinty --> vertical asymp. and that y for which x tends to infinty -->horizontal "

OpenStudy (shubhamsrg):

infinity*

OpenStudy (anonymous):

What you say is like definition of "vertical asymp" and "horizontal asymp" right. I am looking for whether there is specific formula to solve it?

OpenStudy (mimi_x3):

\[\frac{x^{2}}{(x-9)^{2}} \] Vertical asymptote: \((x-9)^2=0\)

OpenStudy (anonymous):

so if I want to get vertical asymptote, I just make the denominator equals to zero?

OpenStudy (mimi_x3):

Yes, for the vertical asymptote.

OpenStudy (cwrw238):

when x = 9 denominator = 0 vertical asymptote at x = 9

OpenStudy (anonymous):

how about horizontal one?

OpenStudy (shubhamsrg):

find x in terms of y and see where does x tends to infinty

OpenStudy (anonymous):

you mean ,use y to express x? y=x^2 over (x-9)^2 then y(x-9)^2=x^2 then... x=sth *y ,right?

OpenStudy (shubhamsrg):

you can take root of both sides..that'll give you sqrt(y) and -sqrt(y) and it'll help you get x in a much simplified form

OpenStudy (shubhamsrg):

sqrt i mean by root**

OpenStudy (anonymous):

you mean dy/dx*y ' = [x^2 over (x-9)^2] ' ? derivative both sides, right??? haha... English is my second language.. thanks for your patience!

OpenStudy (mimi_x3):

You don't differentiate to find the horizontal asympotes :/

OpenStudy (anonymous):

i misunderstood ?

OpenStudy (shubhamsrg):

i mean taking square root of both sides.. like for eg. if x^2 = 4 then x=2 or -2 you take sqrt here..was it clearer mam ?

OpenStudy (anonymous):

ohhh ! yep

OpenStudy (cwrw238):

x / x-9 : as x approaches negative infinity y approaches 0

OpenStudy (anonymous):

I found a way to find horizontal asymptotes!!!! 1) Put equation or function in standard form. 2) Remove everything except the biggest exponents of x found in the numerator and denominator.

OpenStudy (anonymous):

thanksssssssss you guyyyyyssss <3 <3 <3

OpenStudy (shubhamsrg):

hmm..

OpenStudy (anonymous):

haha